Suppose movie downloads cost $1 apiece and game downloads cost $3. If the marginal utility of movie downloads at the optimal mix of consumption is 10 utils, what is the marginal utility of a game download?

Instructions: Enter your answer as a whole number.

utils
Answer #1

Answer

30 utils

The optimum consumption is where per dollar marginal utility of both goods is equal.

marginal utility of movie downloads /cost of movie downloads = marginal utility of game downloads/ cost of game downloads

10/1= marginal utility of game downloads/3

marginal utility of game downloads=10*3=30 utils

the marginal utility of game downloads is 30 utils at the optimal mix of consumption level.